the base of a triangle is 3 times as long as its height. if the area of the triangle is 24 cm squared, find the length of the base and the height.
Answer by venugopalramana(3286)

HT =H
BASE IS 3 TIMES =3H
AREA =BH/2=24
3H*H/2=24
H*H=24*2/3=16
H=4